A stock just paid a dividend of $1.97. The dividend is expected to grow at 20.26% for three years and then grow at 3.71% thereafter. The required return on the stock is 13.75%. What is the value of the stock?
Answer #1

D1=(1.97*1.2026)=2.369122

D2=(2.369122*1.2026)=2.849106117

D3=(2.849106117*1.2026)=3.426335016

Value after year 3=(D3*Growth Rate)/(Required rate-Growth rate)

=(3.426335016*1.0371)/(0.1375-0.0371)

=$35.39294866

Hence current value=Future dividends and value*Present value of discounting factor(rate%,time period)

=2.369122/1.1375+2.849106117/1.1375^2+3.426335016/1.1375^3+35.39294866/1.1375^3
=$30.66(Approx).